Showing all working, find a formula for the general term of the sequence (sn) = s3, s4, s5 . . . defined by

s3 = π and sn = sn-1 - 2 for n >= 4.


1
Expert's answer
2021-12-07T13:56:40-0500

An arithmetic progression is a sequence of the form


"a, a + d, a + 2d, \u2026 , a + nd, \u2026"

where the initial term "a" and the common difference "d" are real numbers.

We have

"a=s_3=\pi,"

"d=s_n-s_{n-1}=s_{n-1}-2-s_{n-1}=-2, n\geq4"

Hence we have the arithmetic progression "\{s_n\}"


"\pi, \pi+(-2), \pi+2(-2), \pi+3(-2),...,\pi+(n-3)(-2),..."

The general term of the sequence is


"s_n=\pi-2(n-3), n\ge 3"